OpenStudy (stormoshawty):

Which is a solution of the equation y=5x+3? A. (5, 28) B. (28, 5) C. (4, 20) D. (3, 19)

OpenStudy (anonymous):

D.(3,19)

OpenStudy (shamil98):

wrong.

OpenStudy (shamil98):

28 = 5(5) + 3 28 = 28 it's A

OpenStudy (anonymous):

A) y= (5)(5) +3 = 28

OpenStudy (monkey202):

yeah its a lol

OpenStudy (anonymous):

ops sorry, @Stormoshawty i thought the q' was asking the roots but im sorry i just realised that it wouldnt make sense in that case and yes @shamil98 is right

OpenStudy (anonymous):

5x-y+3=0 or 5x-y=-3. You could also multiply both sides of any equation by the same non-zero number to get an equivalent expression. In particular x/(3/5) - y/(3) = -1 the x-intercept and y-intercept dividing x and y respectively. (5,28)